# 37N35, 65F55, 90C22, 90C59, 93B52

### v1

@@ Abstract
We introduce two multi-fidelity variants an adaptive sampling scheme for the data-driven construction of stabilizing feedback controllers for nonlinear dynamical systems. The proposed method generates informative data while simultaneously stabilizing the system to avoid instabilities during the data collection, which otherwise could render the collected data meaningless.
@@ 2.1 Stabilization of dynamical systems
We are interested in stabilizing processes towards their steady states. A steady state $(\bar{x},\bar{u})$ is an equilibrium of the time evolution function such that in the continuous-time case eq. 1 it holds
@@ 2.3 Data informativity and controller inference
The deviation from the desired steady state $x(t)=x^{\operatorname{n}}(t)-\bar{x}$ is then approximated by the linear systems
@@
Let $(U,X,Y)$ be a data triplet sampled from a linear state-space model of dimension $N$ and let $\mathcal{V}$ be the $r$-dimensional subspace spanned by all trajectories of the model such that for an orthogonal basis matrix $V$ of $\mathcal{V}$, the state satisfies $x(t)=V\hat{x}(t)$ for all $t\geq 0$, where $\hat{x}(t)\in\mathbb{R}^{r}$ is the state of a linear $r$-dimensional model. The data $(U,X,Y)$ are informative for stabilization over $\mathcal{V}$ if and only if $(U,\widehat{X},\widehat{Y})$ satisfies Proposition 1, with
@@
A stabilizing controller is then given by $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$, where $\widehat{K}$ is a controller that stabilizes all linear models that describe the projected data $(U,\widehat{X},\widehat{Y})$.
Proposition 2 allows to design a stabilizing controller $K$ for an $N$-dimensional model using only $r$ many data samples. The state equality in Proposition 2 is necessary to give a guarantee for the stabilization property; however, in practice it has been observed that already an approximate relation $x(t)\approx V\hat{x}(t)$ for some dimension $r<n_{\min}$, where $n_{\min}$ is the intrinsic system dimension, is sufficient to design a stabilizing controller via Proposition 2. The application of the stabilization theory to nonlinear systems is possible via system linearization based on eq. 7. Under the assumption that the model $f$ is analytic, it holds that for $x^{\operatorname{n}}(t)$ staying close enough to the steady state $\bar{x}$, a controller that stabilizes the corresponding linearization eq. 8 also stabilizes the nonlinear system; see [42, Prop. 3.1]. With a similar argument, we can use the shifted data $(U,\widetilde{X},\widetilde{Y})$ as approximately linear data because the relation eq. 10 holds and design a stabilizing controller from Propositions 1 and 2. The resulting method for low-dimensional controller inference of nonlinear systems is summarized in Algorithm 1. The algorithm operates on the data from the nonlinear model eq. 1 or eq. 2. We call the controller inferred by Algorithm 1 stabilizing over the subspace $\mathcal{V}$ with basis $V$ as it is stabilizing all nonlinear systems with linearizations that describe the shifted and projected data
@@
4Compute the reduced and shifted data triplet $(U,\widehat{X},\widehat{Y})$ via
@@
6Infer a low-dimensional stabilizing feedback $\widehat{K}=U\Theta(\widehat{X}\Theta)^{-1}$ for $(U,\widehat{X},\widehat{Y})$ solving either eq. 11 or eq. 12 for the unknown $\Theta$.
7
8Lift the inferred controller $\widehat{K}$ to the high-dimensional space via $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$.
@@ 3.1 Iterative controller inference
which locally stabilize the system eq. 1 or eq. 2 over a sequence of nested subspaces with non-decreasing dimensions $r_{1}\leq r_{2}\leq\dots\leq r_{T}$,
@@
$U^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}u^{\operatorname{n}}(t_{0})&u^{\operatorname{n}}(t_{1})&\ldots&u^{\operatorname{n}}(t_{j-1})\end{bmatrix},$
$X^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}x^{\operatorname{n}}(t_{0})&x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{j-1})\end{bmatrix},$
$Y^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}f(x^{\operatorname{n}}(t_{0}),u^{\operatorname{n}}(t_{0}))&f(x^{\operatorname{n}}(t_{1}),u^{\operatorname{n}}(t_{1}))&\ldots&f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}(t_{j-1}))\end{bmatrix},$
to the current data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$. To extend the data triplet, we generate an input signal $u(t_{j})$ as
@@
$y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}(t_{j-1})).$ (14)
@@
$\mathcal{V}_{j}=\mathcal{V}_{j-1}\cup\operatorname{span}(x^{\operatorname{n}}(t_{j-1})-\bar{x}).$
To advance the state to $x^{\operatorname{n}}(t_{j})$, we integrate the system in time using $u^{\operatorname{n}}_{j-1}(t)$ from eq. 13. Note that in the case of discrete-time systems eq. 2 and $t_{j}=t_{j-1}+1$, the system query and integration over time are equivalent such that $y^{\operatorname{n}}(t_{j-1})=x^{\operatorname{n}}(t_{j})$. In the first iteration $j=1$, the initial matrices $U^{\operatorname{n}}_{0}$, $X^{\operatorname{n}}_{0}$, and $Y^{\operatorname{n}}_{0}$ are empty, and there is no controller $K_{0}$ available. Therefore, we use a normally distributed random input signal $u(t)$ to query the system at time $t_{0}$ and to integrate the state from $x^{\operatorname{n}}(t_{0})$ to $x^{\operatorname{n}}(t_{1})$. We discuss this in more in detail in Section 3.4.
@@ 3.2 Adaptive inputs for sampling informative data
Since $f$ is describes a linear model, there exist matrices $A\in\mathbb{R}^{N\times N}$ and $B\times\mathbb{R}^{N\times p}$ such that
@@
with $\widehat{A}\in\mathbb{R}^{n_{\min}\times n_{\min}}$ and $\widehat{B}\in\mathbb{R}^{n_{\min}\times p}$. The system matrices in eq. 17 are obtained via projection onto $\mathcal{V}_{\min}$ using the orthogonal basis matrix $V$ such that $\widehat{A}=V^{\mkern-1.5mu\mathsf{T}}AV$ and $\widehat{B}=V^{\mkern-1.5mu\mathsf{T}}B$. Via Proposition 2, we know that $(U,X,Y)$ is informative for stabilization over $\mathcal{V}_{\min}$ if the projected data set $(U,\widehat{X},\widehat{Y})$, where $X=V\widehat{X}$ and $Y=V\widehat{Y}$, is informative for stabilization. Inserting the projected data into eq. 16 yields
@@
is stable, too. From eq. 19, we then get that the product $\widehat{Y}\widehat{X}^{\dagger}$ is stable. By applying part (a) of Proposition 1, the projected data triplet $(U,\widehat{X},\widehat{Y})$ is informative for stabilization such that from Proposition 2, the result of this lemma follows. ∎
@@
Let $f$ be a model with unstable steady state $(\bar{x},\bar{u})$ from which the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ has been sampled, with the state data $X^{\operatorname{n}}=\begin{bmatrix}x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{T})\end{bmatrix}\in\mathbb{R}^{N\times T}$ being close to the steady state such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ holds for some $\epsilon>0$ small enough and $j=1,\ldots,T$. Also, let the the shifted data $\widetilde{X}$ obtained from $X^{\operatorname{n}}$ via eq. 9 have the row rank $n_{\min}$. If $U^{\operatorname{n}}$ is chosen via eq. 5 for a $K$ that stabilizes the system described by $f$ locally, then the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization by linear feedback.
@@
First, we observe that for $\epsilon$ small enough, the shifted data triplet $(U,\widetilde{X},\widetilde{Y})$ obtained via eq. 9 approximates the idealized data $(U,X,Y)$ of the linearization eq. 8 arbitrarily well such that following the theory in [23, Sec. 10.1], we may assume that the data $(U,\widetilde{X},\widetilde{Y})$ can be described by the linearization eq. 8. Then, the $n_{\min}$-dimensional subspace $\mathcal{V}_{\min}$ of all trajectories of the linearization is spanned by the orthogonal basis matrix resulting from a singular value decomposition of the given data since $\widetilde{X}$ is assumed to have row rank $n_{\min}$ such that
$\begin{bmatrix}V_{\min}&V_{2}\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&0\end{bmatrix}W=X,$
where the basis matrix $V_{\min}$ corresponds to the $n_{\min}$ nonzero singular values in $\Sigma$, and $W$ contains the right singular vectors. With the choice of $U^{\operatorname{n}}$ via eq. 5 with a locally stabilizing controller $K$, the assumptions of Lemma 2 are satisfied for the linearization eq. 8, which means that the shifted data triplet $(U,\widetilde{X},\widetilde{Y})$ is informative for stabilization over $\mathcal{V}_{\min}$. From [42, Cor. 1] it follows that if $(U,\widetilde{X},\widetilde{Y})$ is informative for stabilization, then $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization over $\mathcal{V}_{\min}$, which concludes the proof. ∎
@@ 3.3 Convergence analysis
Let the iteratively constructed data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ defined in eq. 15 from the ICI method be such that the concatenated matrix $\begin{bmatrix}\widetilde{X}_{j}^{\mkern-1.5mu\mathsf{T}}&U_{j}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ has full column rank for all $j=1,\ldots,n_{\min}+p$, where $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ and $U_{j}\in\mathbb{R}^{p\times j}$ are shifted data matrices eq. 9, and the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough. Then, it is guaranteed that a stabilizing controller is inferred with ICI from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ for $j=n_{\min}+p$.
@@
Let $\mathcal{V}_{j}\subseteq\mathbb{R}^{N}$ be a $j$-dimensional subspace and let $K_{j-1}\in\mathbb{R}^{p\times N}$ be a state-feedback controller that is stabilizing the system described by the model $f$ over the subspace $\mathcal{V}_{j}$. Given the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ from the ICI method as in eq. 15, we assume that the shifted state samples $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ obtained from $X^{\operatorname{n}}_{j}$ via eq. 9 have full column rank $j$, the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ in the sense that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough, and the input samples $U^{\operatorname{n}}_{j}$ satisfy eq. 5 columnwise with the state samples $X^{\operatorname{n}}_{j}$ and the controller $K_{j-1}$. Then, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is informative for stabilization over $\mathcal{V}_{j}$. If $j=n_{\min}$ and $\mathcal{V}_{j}$ is so that $x^{\operatorname{n}}(t)-\bar{x}\in\mathcal{V}_{j}$ for all $x^{\operatorname{n}}(t)$ generated by the model $f$ with $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$, then the controller inferred from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is guaranteed to locally stabilize the underlying system.
@@
From the construction of the input samples $U^{\operatorname{n}}_{j}$ and the full column rank of the state samples $X^{\operatorname{n}}_{j}$, the conditions of Theorem 1 are satisfied for the nonlinear system projected onto the low-dimensional subspace $\mathcal{V}_{j}$. Therefore, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ must be informative for stabilization over $\mathcal{V}_{j}$. Since the controller inferred from the informative data set is stabilizing all linearizations that describe the shifted and projected data, it follows that if $j=n_{\min}$ and $\mathcal{V}_{j}$ containing all state trajectories generated by $f$, the linearization of the underlying system is stabilized by the inferred controller. Therefore, the controller is guaranteed to locally stabilize the underlying system, which concludes the proof. ∎
@@
Considering the assumptions made in Theorem 2, we have to note that in the ICI method as presented in Section 3.1, the assumption that the same $K_{j-1}$ has been used for the generation of $U^{\operatorname{n}}_{j}$ will not be satisfied since the controller is updated in every step of the method. However, we provide a numerical extension of the method to satisfy this assumption with further details in Section 3.5. Also, we can see that the assumption of $K_{j-1}$ being stabilizing over $\mathcal{V}_{j}$ can be interpreted as $\mathcal{V}_{j}$ not containing more unstable dynamics than $\mathcal{V}_{j-1}$. To see this, we use from Section 3.1 the fact that in ICI, we have that $\mathcal{V}_{j-1}\subseteq\mathcal{V}_{j}$ holds for all $j\geq 1$, and we assume for simplicity of presentation that the corresponding basis matrices $V_{j-1}$ and $V_{j}$ are computed via the Gram-Schmidt procedure so that the columns of $V_{j-1}$ are the first $r_{j-1}$ many columns of $V_{j}$. For the controller obtained in the previous iteration via $K_{j-1}=\widehat{K}_{j-1}V_{j-1}^{\mkern-1.5mu\mathsf{T}}$, it follows that for all possible states from the extended subspace $x\in\mathcal{V}_{j}$ it holds
@@
This shows that the controller $K_{j-1}$ takes the same actions for states from $\mathcal{V}_{j}$ as for states from $\mathcal{V}_{j-1}$, and if $\mathcal{V}_{j}$ does contain unstable dynamics that are not contained in $\mathcal{V}_{j-1}$, then $K_{j-1}$ cannot be stabilizing. Note that if $K_{j-1}$ is not stabilizing over $\mathcal{V}_{j}$, this does not imply that there is no $K_{j}$ stabilizing over $\mathcal{V}_{j}$ or that the corresponding data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is not informative for stabilization; see the discussions in [43].

### v2

@@ Abstract
We introduce an adaptive sampling scheme for the data-driven construction of stabilizing feedback controllers for nonlinear dynamical systems. The proposed method generates informative data while simultaneously stabilizing the system to avoid instabilities during the data collection, which otherwise could render the collected data meaningless.
@@ 2.1 Stabilization of dynamical systems
The superscript $\operatorname{n}$ used above in the notation of the states $x^{\operatorname{n}}(t)$ and inputs $u^{\operatorname{n}}(t)$ indicates throughout this work the association of these quantities with the original, potentially nonlinear system described by the model $f$. We are interested in stabilizing processes towards their steady states. A steady state $(\bar{x},\bar{u})$ is an equilibrium of the time evolution function such that in the continuous-time case eq. 1 it holds
@@ 2.3 Data informativity and controller inference
The deviation from the desired steady state $\tilde{x}(t)=x^{\operatorname{n}}(t)-\bar{x}$ is then approximated by the state $x(t)$ of the linear systems
@@
[…]
@@
[…]
@@
[…]
@@
6-
[…]
8
9Lift the inferred controller $\widehat{K}$ to the high-dimensional space via $K=\widehat{K}V^{\mkern-1.5mu\mathsf{T}}$.
@@ 3.1 Iterative controller inference
which in lifted form locally stabilize the system eq. 1 or eq. 2 over a sequence of nested subspaces with non-decreasing dimensions $r_{1}\leq r_{2}\leq\dots\leq r_{T}$,
@@
$U^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}u^{\operatorname{n}}_{0}(t_{0})&u^{\operatorname{n}}_{1}(t_{1})&\ldots&u^{\operatorname{n}}_{j-2}(t_{j-2})\end{bmatrix}$,
$X^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}x^{\operatorname{n}}(t_{0})&x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{j-2})\end{bmatrix}$,
$Y^{\operatorname{n}}_{j-1}$ $=\begin{bmatrix}f(x^{\operatorname{n}}(t_{0}),u^{\operatorname{n}}_{0}(t_{0}))&f(x^{\operatorname{n}}(t_{1}),u^{\operatorname{n}}_{1}(t_{1}))&\ldots&f(x^{\operatorname{n}}(t_{j-2}),u^{\operatorname{n}}_{j-2}(t_{j-2}))\end{bmatrix}$,
to the current data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$. To extend the data triplet, we generate an input signal $u(t)$ as
@@
$y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}_{j-1}(t_{j-1})).$ (14)
@@
$\mathcal{V}_{j}=\mathcal{V}_{j-1}\cup\operatorname{span}(x^{\operatorname{n}}(t_{j-1})-\bar{x}),$
except in certain special cases, which are further discussed in Section 3.4. To advance the state to $x^{\operatorname{n}}(t_{j})$, we integrate the system in time using $u^{\operatorname{n}}_{j-1}(t)$ from eq. 13. Note that in the case of discrete-time systems eq. 2 and if the discrete time points for the data sampling are only one time step apart so that $t_{j}=t_{j-1}+1$ holds, the system query and integration over time are equivalent such that $y^{\operatorname{n}}(t_{j-1})=x^{\operatorname{n}}(t_{j})$. In the first iteration $j=1$, the initial matrices $U^{\operatorname{n}}_{0}$, $X^{\operatorname{n}}_{0}$, and $Y^{\operatorname{n}}_{0}$ are empty, and there is no controller $K_{0}$ available. Therefore, we use a normally distributed random input signal $u(t)$ to query the system at time $t_{0}$ and to integrate the state from $x^{\operatorname{n}}(t_{0})$ to $x^{\operatorname{n}}(t_{1})$. We discuss this in more in detail in Section 3.4.
@@
This iterative process can be concluded when the dimension of the constructed subspace $\mathcal{V}_{T}$ reaches the intrinsic system dimension $n_{\min}$. Typically, the intrinsic dimension $n_{\min}$ is unknown for general systems. However, stagnation in the expansion of the subspace $\mathcal{V}_{T}$ indicates that no further independent data can be generated via simulation and that $\mathcal{V}_{T}$ has reached its maximum dimension $n_{\min}$. For the practical use of the method in scarce data applications, we propose to end the iteration scheme via a criterion of the form $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert<\operatorname{tol}$ with a user-given tolerance $\operatorname{tol}>0$, which stops the iterations when the system trajectories are close enough to the steady state and can be considered as stabilized.
@@ 3.2 Adaptive inputs for sampling informative data
Since $f$ describes a linear model, there exist matrices $A\in\mathbb{R}^{N\times N}$ and $B\times\mathbb{R}^{N\times p}$ such that
@@
[…]
@@
[…]
@@
Let $f$ be a model with unstable steady state $(\bar{x},\bar{u})$ from which the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ has been sampled, with the state data $X^{\operatorname{n}}=\begin{bmatrix}x^{\operatorname{n}}(t_{1})&\ldots&x^{\operatorname{n}}(t_{T})\end{bmatrix}\in\mathbb{R}^{N\times T}$ being close to the steady state such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ holds for some $\epsilon>0$ small enough and $j=1,\ldots,T$. Also, let the the shifted data $\widetilde{X}$ obtained from $X^{\operatorname{n}}$ via eq. 9 have the row rank $n_{\min}$, the intrinsic dimension of the system. If $U^{\operatorname{n}}$ is chosen via eq. 5 for a $K$ that stabilizes the linearization of $f$, then the data $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization by linear feedback.
@@
[…]
$\begin{bmatrix}V_{\min}&V_{2}\end{bmatrix}\begin{bmatrix}\Sigma&0\\ 0&0\end{bmatrix}W=\widetilde{X}$,
[…]
@@ 3.3 Convergence analysis
Let the iteratively constructed data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ defined in eq. 15 from the ICI method be such that the concatenated matrix $\begin{bmatrix}\widetilde{X}_{j}^{\mkern-1.5mu\mathsf{T}}&U_{j}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}$ has full column rank for all $j=1,\ldots,n_{\min}+p$, where $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ and $U_{j}\in\mathbb{R}^{p\times j}$ are shifted data matrices eq. 9, and the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ such that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough. Then, the controller inferred by ICI from the data $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ for $j=n_{\min}+p$ is guaranteed to stabilize the system described by the model from which the data has been sampled.
@@
Let $\mathcal{V}_{j}\subseteq\mathbb{R}^{N}$ be a $j$-dimensional subspace and let $K_{j-1}\in\mathbb{R}^{p\times N}$ be a state-feedback controller that locally stabilizes the projected system described by the model $V_{j}V_{j}^{\mkern-1.5mu\mathsf{T}}f$, for an orthogonal basis $V_{j}$ of the subspace $\mathcal{V}_{j}$. Given the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ from the ICI method as in eq. 15, we assume that the shifted state samples $\widetilde{X}_{j}\in\mathbb{R}^{N\times j}$ obtained from $X^{\operatorname{n}}_{j}$ via eq. 9 have full column rank $j$, the columns of $X^{\operatorname{n}}_{j}$ are sufficiently close to the steady state $\bar{x}$ in the sense that $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$ for $\epsilon>0$ small enough, and the input samples $U^{\operatorname{n}}_{j}$ satisfy eq. 5 columnwise with the state samples $X^{\operatorname{n}}_{j}$ and the controller $K_{j-1}$. Then, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is informative for stabilization over $\mathcal{V}_{j}$. If $j=n_{\min}$ and $\mathcal{V}_{j}$ is so that $x^{\operatorname{n}}(t)-\bar{x}\in\mathcal{V}_{j}$ for all $x^{\operatorname{n}}(t)$ generated by the model $f$ with $\lVert x^{\operatorname{n}}(t_{j})-\bar{x}\rVert\leq\epsilon$, then the controller inferred from $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is guaranteed to locally stabilize the underlying system.
@@
From the construction of the input samples $U^{\operatorname{n}}_{j}$ and the full column rank of the state samples $X^{\operatorname{n}}_{j}$, the conditions of Theorem 1 are satisfied for the nonlinear system projected onto the low-dimensional subspace $\mathcal{V}_{j}$, which is described by the model $V_{j}V_{j}^{\mkern-1.5mu\mathsf{T}}f(x(t),u(t))$. Therefore, the data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ must be informative for stabilization over $\mathcal{V}_{j}$. Since the controller inferred from the informative data set is stabilizing all linearizations that describe the shifted and projected data, it follows that if $j=n_{\min}$ and $\mathcal{V}_{j}$ containing all state trajectories generated by $f$, the linearization of the underlying system is stabilized by the inferred controller. Therefore, the controller is guaranteed to locally stabilize the underlying system, which concludes the proof. ∎
@@
We note that Theorem 2 critically relies on the assumption that $K_{j-1}$ stabilizes over the subspace $\mathcal{V}_{j}$. If this assumption does not hold, then our ICI method is still applicable but it is not guaranteed to give a minimal data set. Furthermore, considering the assumptions made in Theorem 2, we have to note that in the ICI method as presented in Section 3.1, the assumption that the same $K_{j-1}$ has been used for the generation of $U^{\operatorname{n}}_{j}$ will not be satisfied since the controller is updated in every step of the method. However, we provide a numerical extension of the method to satisfy this assumption with further details in Section 3.5. Also, we can see that the assumption of $K_{j-1}$ being stabilizing over $\mathcal{V}_{j}$ can be interpreted as $\mathcal{V}_{j}$ not containing more unstable dynamics than $\mathcal{V}_{j-1}$. To see this, we use from Section 3.1 the fact that in ICI, we have that $\mathcal{V}_{j-1}\subseteq\mathcal{V}_{j}$ holds for all $j\geq 1$, and we assume for simplicity of presentation that the corresponding basis matrices $V_{j-1}$ and $V_{j}$ are computed via the Gram-Schmidt procedure so that the columns of $V_{j-1}$ are the first $r_{j-1}$ many columns of $V_{j}$. For the controller obtained in the previous iteration via $K_{j-1}=\widehat{K}_{j-1}V_{j-1}^{\mkern-1.5mu\mathsf{T}}$, it follows that for all possible states from the extended subspace $x\in\mathcal{V}_{j}$ it holds
@@
This shows that the controller $K_{j-1}$ takes the same actions for states from $\mathcal{V}_{j}$ as for states from $\mathcal{V}_{j-1}$, and if $\mathcal{V}_{j}$ does contain unstable dynamics that are not contained in $\mathcal{V}_{j-1}$, then $K_{j-1}$ cannot be stabilizing. Note that if $K_{j-1}$ is not stabilizing over $\mathcal{V}_{j}$, this does not imply that there is no $K_{j}$ stabilizing over $\mathcal{V}_{j}$ or that the corresponding data triplet $(U^{\operatorname{n}}_{j},X^{\operatorname{n}}_{j},Y^{\operatorname{n}}_{j})$ is not informative for stabilization; see the discussions in [43]. We discuss the handling of intermediate non-informative data sets are in Section 3.4.